OpenStudy (kenkaneki12280000):

What is the solution to the system? Line 1 x y –2 –4 –1 –1 Line 2 x y 2 3 3 1 ( , )

OpenStudy (kenkaneki12280000):

@alivejeremy @Hayhayz

OpenStudy (jameshorton):

The slope is found by y_{2} -y_{1} / x_{2} - x_{1} Line 1: y2 is 6, y1 is 4, x2 is 2, x1 is -2 6-4/2- -2= 2/4=1/2 Slope intercept is written as y= mx+b. m is the slope, and b is the intercept (the intercept is the value of y when x is 0). To find out the intercept, we plug in values of y, m and x 4= 1/2(-2) +b 4=-1+b 5=b The equation for line 1 is y= 1/2x +5 Using the same formulas as above, the equation for line 2 is y= 3x+4 The solution to the system is a point where both lines would meet in the graph, or a set of x and y values that satisfies both equations. y= 3x+4 y= 1/2x+5 3x+4 = 1/2x+5 3x= 1/2x +5 -4 3x= 1/2x +1 3x-1/2x= 1 2.5x=1 2.5x/2.5=1/2.5 x=0.4 Plug 0.4 into the slope intercept equation

OpenStudy (jameshorton):

y= 3(0.4) +4 y= 1.2 +4 y=5.2 y= 1/2(0.4) +5 y= 0.2 +5 y= 5.2 The solution is 0.4, 5.2 The other set of lines Line 1: y= 3x+2 Line 2: y= -2x +7 Solution: 1,5

OpenStudy (kenkaneki12280000):

thx ill tell u if u are correct altho im pretty sure u are ;)

OpenStudy (kenkaneki12280000):

your answer was right

OpenStudy (jameshorton):

medal me plz

OpenStudy (kenkaneki12280000):

ill medal u dont wrry but can u help meh with 5 more if u have the time to spare

OpenStudy (jameshorton):

ok

OpenStudy (kenkaneki12280000):

ok first question is The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –3 5 0 4 Line 2 x y 0 4 6 2 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (jameshorton):

okay. so, first you need the slope of both lines. Slope equation is (y2-y1)/(x2-x1) (Line 1) 2--2/1--1 4/2 2 (Line 2)

OpenStudy (jameshorton):

(1-10)/(4--2) -9/6 -3/2

OpenStudy (kenkaneki12280000):

does that mean its a b c or d

OpenStudy (kenkaneki12280000):

the second question is The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –4 5 0 2 Line 2 x y 0 5 4 2 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (jameshorton):

it b

OpenStudy (kenkaneki12280000):

for the first or second

OpenStudy (jameshorton):

first

OpenStudy (kenkaneki12280000):

and the second when ever u are ready

OpenStudy (kenkaneki12280000):

ill just post it on a diff one brb and then ill tag

OpenStudy (jameshorton):

ok